# Multivector Reranking in the Era of Strong First-Stage Retrievers

### v2

@@ 4.1 Reproducibility and Comparison with the State of the Art
Table 3: Evaluation of reranking optimizations. End-to-end (1st + 2nd stage) retrieval time (ms) on MsMarco.
MRR@10
0.400 0.402 0.404 0.405
Half-precision $2.5$ $3.7$ $6.7$ -
   + CP $2.2$ $3.2$ $5.7$ $9.8$
   + EE $2.2$ $3.2$ $5.6$ -
   + CP + EE $2.1$ $2.8$ $5.6$ -
@@ 4.2 Impact of Quantization and Reranking Optimizations
We perform an ablation study on the MsMarco dataset to analyze the effect of multivector quantization and our proposed reranking optimizations, i.e. Candidate Pruning (CP) and Early Exit (EE). Table 3 and Figure 2 report MRR@10 versus reranking latency for all tested configurations on the MsMarco dataset, where on first stage we employ Splade CoCondenser representations.
Table 3 isolates the Half-precision configuration. This setup defines the upper bound in retrieval quality, reaching MRR@10 above $0.40$, but at a very high memory cost ($256$ Bytes per token embedding). Latency ranges between 2.5 and 9.8 ms per query, depending on the MRR target. Both CP and EE reduce latency by 10–20% with no impact on quality, while their combination (CP+EE) offers the best balance, achieving up to $2.1$ ms per query at similar MRR.
The plots in Figure 2 compare all five compression schemes, i.e., Half-precision, OPQ64, MOPQ32, JMPQ32 and JMPQ16, under different optimization settings. All compressed variants exhibit slightly lower MRR (up to $\approx$0.399) but deliver substantial memory savings, from $4\times$ (OPQ64) to $12.8\times$ (JMPQ16) relative to the Half-precision baseline.
@@
OPQ64 maintains strong accuracy but suffers from the highest latency (up to 13.5 ms) due to the computation of 64 distance tables per query token. MOPQ32 achieves similar performance at roughly half the memory footprint (36 Bytes per embedding), although its peak effectiveness is slightly lower. JMPQ32 and JMPQ16 deliver the best overall trade-offs: they retain most of the effectiveness while improving efficiency and compressing memory by $7\times$$12.8\times$. Across all quantization settings, CP and EE consistently improve efficiency with no losses in MRR. The best results, as in the Table 3, are obtained combining CP and EE, achieving a maximum speed-up of $1.8\times$.

### v3

@@ 4.2 Impact of Quantization and Reranking Optimizations
We perform an ablation study on the MsMarco dataset to analyze the effect of multivector quantization and our proposed reranking optimizations, i.e. Candidate Pruning (CP) and Early Exit (EE). Figure 2 reports MRR@10 versus search latency for all tested configurations on the MsMarco dataset, where on first stage we employ Splade CoCondenser representations.
The figure compare all five compression schemes, i.e., Half-precision, OPQ64, MOPQ32, JMPQ32 and JMPQ16, under different optimization settings. JMPQ32 shows no loss in quality despite a $7\times$ memory reduction compared to the Half-precision setting. The other compressed variants exhibit slightly lower MRR but deliver substantial memory savings, ranging from $4\times$ (OPQ64) to $12.8\times$ (JMPQ16) relative to the baseline.
@@
OPQ64 maintains strong accuracy but suffers from the highest latency (up to 13.5 ms) due to the computation of 64 distance tables per query token. MOPQ32 achieves similar performance at roughly half the memory footprint (36 Bytes per embedding), although its peak effectiveness is slightly lower. JMPQ32 and JMPQ16 deliver the best overall trade-offs: they retain most of the effectiveness while improving efficiency and compressing memory by $7\times$$12.8\times$.
Regarding the reranking optimizations, CP and EE consistently improve efficiency across all quantization settings, achieving a maximum speed-up of $1.8\times$. CP is the most reliable optimization; by relying on uncompressed first-stage scores, it consistently reaches the same MRR as its unoptimized counterpart. Conversely, EE is occasionally too aggressive, preventing the algorithm from reaching top accuracy; this also affects the combined EE + CP setting.